An aptitude test is designed to measure leadership abilities of the test subjects. Suppose that the scores on the test are normally distributed with a mean of 550 and a standard deviation of 125. The individuals who exceed 700 on this test are considered to be potential leaders. What proportion of the population are considered to be potential leaders? Round your answer to at least four decimal places.
Answer #1

Solution :

Given that ,

mean = = 550

standard deviation = = 125

P(x > 700) = 1 - P(x < 700)

= 1 - P((x - ) / < (700 - 550) / 125)

= 1 - P(z < 1.2)

= 1 - 0.8849   

= 0.1151

considered to be potential leaders = 0.1151